OpenStudy (anonymous):

write -4 4/9 as a improper fraction

OpenStudy (anonymous):

numerator is -4*9-4=-40, so -40/9

OpenStudy (anonymous):

another way: write the -4 with a denominator of 9 and then subtract 4/9:\[\frac{-4}{1}*\frac{9}{9}-\frac{4}{9}=\frac{-36-4}{9}=\frac{-40}{9}\]